OpenStudy (anonymous):

What is 3 root of 135? A. 5 3 root of 2 B. 2 3 root of 5 C. 5 3 root of 3 D. 3 3 root of 5

OpenStudy (anonymous):

which one do you think is you answer love?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

what do you think d?

OpenStudy (michele_laino):

hint: we can write this: \[\sqrt {135} = \sqrt {5 \times {3^2} \times 3} \]

OpenStudy (anonymous):

b

OpenStudy (michele_laino):

if we consider the third-root, then we have: \[\sqrt[3]{{135}} = \sqrt[3]{{5 \times {3^3}}} = 3\sqrt[3]{5}\]

OpenStudy (anonymous):

so-is-it-d

OpenStudy (michele_laino):

\[\Large \sqrt[3]{{135}} = \sqrt[3]{{5 \times {3^3}}} = 3\sqrt[3]{5}\]

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

I confirm it is the option d.

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

can-u-help-with-my-next-1

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

its-the-same

OpenStudy (michele_laino):

namely?